OpenStudy (anonymous):

I got a couple more questions that i am confused about! Question 1: Describe how you would use the distributive property to simplify (39 x 5). (I think it may be 195 but i dont know if i worked it out right) Question 2: Your best friend tells you the solution of 375 = x - 28 is 347. Explain your best friend's error and tell him/her the correct answer

OpenStudy (anonymous):

for number 2 375=x-28 Add 28 to both sides so it can cancel out X=403 And plug it in and you can check it like that.

OpenStudy (anonymous):

Question 2 = Add 28 on both sides Question 1 = (39*5) = 195 Put 39 into a+b form so that (a+b)5 = a*5 +b*5 = 195 a+b = 39 a*5 +b*5 = 195 solve the equations for a and b

hero (hero):

Or just simply do this: 5(39) = 5(30 + 9)

OpenStudy (anonymous):

@Hero I assumed from the second equation. But I apologize if its not the correct way for Timetolearn

hero (hero):

I see. For questions like these. If you have something like 27 you can rewrite it as 20 + 7

hero (hero):

Notice that "thirty nine" is actually 30 + 9 "twenty seven" is actually 20 + 7

OpenStudy (anonymous):

Oh! Thanks so much. This is some preview questions for a test :) I just did not understand well.

hero (hero):

You're welcome

hero (hero):

Also, @sangya21 you attempted to represent the problem as a system of equations. It wasn't exactly meant to be represented in this manner even though your system works. This approach doesn't work for all systems of equations obviously.

OpenStudy (anonymous):

@Hero It has been a while for me doing these questions and as a course of nature I have forgot few basics. Thanks for clearing it up. I try to solve through the ways I remember